OpenStudy (anonymous):

Each of the letters in the word MATHEMATICS is represented by a letter tile in a bag. Stacey picks three letters, without replacement. To the nearest whole percent, which is the probability that she'll get all vowels?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

total vowels is 4. total letters is 11 P of vowel on first pick = 4/11 P of vowel on 2nd pick= (4/11)( 3/10) (as only 3 vowels left and 10 letters) P of vowel on 3rd pick = (4/11)(3/10)(2/9) (only 2 vowels left and 9 letters) P of vowel on 4th pick = (4/11)(3/10)(2/9)(1/8) (only 1 vowel and 8 letters) so total probability is 24/7920 =0.00303030 recurring=0.3 recurring percent. to nearest whole percent = 0